what is 2 feet more than 1/3 of 3 feet ?
Answer by stanbon(75887)

what is 2 feet more than 1/3 of 3 feet ?
----
Ans (1/3)3+2 = 1+2 = 3
------------------------
Cheers,
Stan H.
--------------